OpenStudy (fanduekisses):

at what values of x do the graphs of y= x^2 and y=-sqrt(x) have perpendicular tangent lines?

OpenStudy (solomonzelman):

have perpendicular tangent lines to each other? (saying that the tangents are perpendicular?)

OpenStudy (solomonzelman):

lets start from finding the derivative of the entire function.

OpenStudy (solomonzelman):

I mean of the 1st function

OpenStudy (jdoe0001):

are you supposed to graph, or just find it on some other way?

OpenStudy (solomonzelman):

other way, jdoe

OpenStudy (jdoe0001):

hmm

OpenStudy (fanduekisses):

y'=2x

OpenStudy (solomonzelman):

We will need the derivative of both of the functions: lemme do a notation rearrangement: \(\normalsize\color{royalblue}{ \rm f(x)= x^2 }\) \(\normalsize\color{royalblue}{ \rm g(x)=-\sqrt{x} }\)

OpenStudy (fanduekisses):

and the second function y'=-1/2(x)^-1/2

OpenStudy (solomonzelman):

yes, \(\normalsize\color{royalblue}{ \rm f'(x)= 2x }\)

OpenStudy (solomonzelman):

and \(\normalsize\color{royalblue}{ \rm g'(x)= \frac{\Large 1 }{\Large2\sqrt{x}} }\)

OpenStudy (solomonzelman):

Note that: \(\normalsize\color{royalblue}{ \rm f'(x)=2x}\) and \(\normalsize\color{royalblue}{ \rm g'(x)= \frac{\Large 1 }{\Large2\sqrt{x}} }\) are the slopes. of both of the functions:

OpenStudy (solomonzelman):

you need to find a value of a, (for x), so that \(\normalsize\color{royalblue}{ \rm g'(a)=~- \frac{\Large 1 }{\Large f'(a)} }\) or, \(\normalsize\color{royalblue}{ \rm f'(a)=~- \frac{\Large 1 }{\Large g'(a)} }\) it is all same

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm f'(x)=~- \frac{\Large 1 }{\Large g'(x)} }\) \(\normalsize\color{royalblue}{ \rm 2x=~- \frac{\Large 1 }{\Large \frac{\Large 1 }{\Large 2\sqrt{x}}} }\) and on...

OpenStudy (solomonzelman):

oh my bad.

OpenStudy (solomonzelman):

you function g(x) is negative so the derivative is negative. \(\normalsize\color{royalblue}{ \rm 2x= \frac{\Large 1 }{\Large \frac{\Large 1 }{\Large 2\sqrt{x}}} }\)

OpenStudy (solomonzelman):

I am setting the perpendicular slope equation, substituting the derivative of the functions into that formula. (2 minus on the left side make it positive)

OpenStudy (fanduekisses):

ok

OpenStudy (fanduekisses):

got lost after you function g(x) is negative so the derivative is negative....

OpenStudy (fanduekisses):

So I look for the derivatives, then set one of them equal to the perpendicular slope of the other one?

OpenStudy (solomonzelman):

You know that: m=-1/m are prep slopes ? (excuse me for these words, mathematicians)

OpenStudy (fanduekisses):

2x= -2sqrt(x)/1

OpenStudy (solomonzelman):

close

OpenStudy (solomonzelman):

you derivative is negative, and you got another negative in the m=-1/m

OpenStudy (solomonzelman):

your deriv.* ...

OpenStudy (solomonzelman):

I mean the derivative for -sqrt{x}

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm f'(x)=-\frac{\Large 1}{\Large g'(x)} }\)

OpenStudy (solomonzelman):

perpendicular slopes, correct?

OpenStudy (fanduekisses):

yes

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm 2x=-\frac{\Large 1}{\Large \frac{\LARGE 1}{\LARGE -2\sqrt{x}}} }\) plugging the derivatives in...

OpenStudy (solomonzelman):

you got it right that time, for everything besides the sign... see why?

OpenStudy (fanduekisses):

wait how is 2sqrt(x) negative?

OpenStudy (solomonzelman):

you had negative sqrt{x} in the begginging, no?

OpenStudy (fanduekisses):

ohh yes

OpenStudy (solomonzelman):

yes, see why you get: \(\normalsize\color{royalblue}{ \rm 2x=-\frac{\Large 1}{\Large \frac{\LARGE 1}{\LARGE -2\sqrt{x}}} }\) >>> \(\normalsize\color{royalblue}{ \rm 2x=2\sqrt{x}}\)

OpenStudy (solomonzelman):

and then the x, is?

OpenStudy (fanduekisses):

sqrt x

OpenStudy (solomonzelman):

?

OpenStudy (solomonzelman):

what do you mean?

OpenStudy (fanduekisses):

\[\sqrt{x}\]

OpenStudy (solomonzelman):

I see, but what do you want to tell me?

OpenStudy (solomonzelman):

All I did so far is that I set the derivatives of f(x) and g(x): ( where: ~ \(\normalsize\color{royalblue}{ \rm f(x)=x^2}\) ~ \(\normalsize\color{royalblue}{ \rm f'(x)=2x}\) ~ \(\normalsize\color{royalblue}{ \rm g(x)=-\sqrt{x}}\) ~ \(\normalsize\color{royalblue}{ \rm g'(x)=-1/(~2\sqrt{x}~)}\)

OpenStudy (solomonzelman):

set the derivatives (the slopes of f(x) and g(x) into a form of: m = -1/m

OpenStudy (solomonzelman):

since I need to find the x, where the slopes of the functions are perpendicular. (and the slopes are the derivatives)

OpenStudy (fanduekisses):

ok -1/1/-2sqrt(x) and -1/2x

OpenStudy (solomonzelman):

no, you need the slopes of the f(x) and g(x) perp to each other: that is why I set: `f'(x) = - 1 / g'(x)` makes sense?

OpenStudy (solomonzelman):

lost, or with me?>

OpenStudy (fanduekisses):

ok so \[2x=\frac{ -1 }{ \frac{ 1 }{ -2\sqrt{x} } }\]

OpenStudy (fanduekisses):

@SolomonZelman

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

and then solve for x.

OpenStudy (solomonzelman):

this x value will be where the slopes of the 2 functions are perpendicular to each other.

OpenStudy (solomonzelman):

(note: exclude extraneous solutions)

OpenStudy (solomonzelman):

lost me, or can solve for x?

OpenStudy (fanduekisses):

I get x= sqrt(x)

OpenStudy (solomonzelman):

yes, that means that x=?

OpenStudy (solomonzelman):

there are 2 solutions, what are they?

OpenStudy (fanduekisses):

\[\pm \sqrt{x}\]

OpenStudy (solomonzelman):

no plus minus

OpenStudy (solomonzelman):

and secondly, x=\sqrt{x} doesn't that imply that x=1 or 0?

OpenStudy (solomonzelman):

0 won't work, because it is not in the g'(x), so we are left with 1.

OpenStudy (solomonzelman):

so the tangents will be perpendicular at x=1.

OpenStudy (solomonzelman):

want to find these tangent lines and graph them to see if it is true?

OpenStudy (fanduekisses):

yes :)

OpenStudy (solomonzelman):

We have: \(\normalsize\color{royalblue}{ f(x)=x^2 }\) \(\normalsize\color{royalblue}{ f'(x)=2x }\) \(\normalsize\color{teal}{ g(x)=-\sqrt{x} }\) \(\normalsize\color{teal}{ g'(x)=-1/(~2\sqrt{x}~) }\)

OpenStudy (solomonzelman):

find the tangent line to f(x), at x=1. (can you do that? )

OpenStudy (solomonzelman):

steps for this: Find \(\normalsize\color{royalblue}{ f'(1) }\) Find \(\normalsize\color{royalblue}{ f(1) }\) (y coordinate of the point) Use the point slope form.

OpenStudy (fanduekisses):

y-1=2(x-1) y=2x-1

OpenStudy (solomonzelman):

yes, we can leave in a point slope form, ain't important

OpenStudy (fanduekisses):

=D

OpenStudy (solomonzelman):

can you find the slope of the tangent line, to the curve g(x), at x=1 ?

OpenStudy (fanduekisses):

y+1=-1/2x-1/2

OpenStudy (solomonzelman):

without expanding, so that I we can clearly see the point plugged: y+1=-1/2(x-1)

OpenStudy (solomonzelman):

since g(1) is the y coordinate of the point, and 1 is the x coordinate: g(1)=-sqrt(1) = -1 the point is (1,-1) . and with a slope -1/2, that gives as: the above line. slope of -1/2 is correct;) f'(1)= -1/(2sqrt(1)) = -1/2

OpenStudy (solomonzelman):

So tangents at \(\normalsize\color{royalblue}{ x=1 }\) For \(\normalsize\color{royalblue}{ f(x) }\): \(\normalsize\color{royalblue}{ y-1=2(x-1) }\) For \(\normalsize\color{teal}{ f(x) }\): \(\normalsize\color{teal}{ y+1=\frac{1}{2}(x-1) }\)

OpenStudy (solomonzelman):

There is the graph of it: https://www.desmos.com/calculator/efm9gxyiam